solving systems of equations. Solve the system: 2x+y=1 and 5x-y=20.
Answer by checkley77(12844)

2x+y=1
5x-y=20 add
--------------------
7x=21
x=21/7
x=3 ans
2*3+y=1
6+y=1`
y=1-6
y=-5 ans.
Proof:
5*3-(-5)=20
15+5=20
20=20